The probability that Kodjo and Adoga pass an examination are 34 and 35 respectively. Find the probability of both boys failing the examination.
1/3
3/10
9/20
1/3
1
Expert's answer
2015-02-25T13:04:12-0500
There is a mistake in the question: there should be 3/4 and 3/5 instead of 34 and 35. The probability of both boys failing the examination is (1-3/4)(1-3/5)=1/4*2/5=1/10 by multiplication rule of probability.